OpenStudy (anonymous):

how do you find 44% of 82

OpenStudy (anonymous):

44% is the same as \[\frac{44}{100}\] (per-cent means per-hundred) which is equal to \[0.44\] So to find 44% of 82 we do:\[0.44 \times 82=36.08\]

OpenStudy (anonymous):

You can just divide 0.44 by 82 and then whatever the answer you get you just move the decimal point 2 places to the left

OpenStudy (anonymous):

You should give me a medal if that helped!!;)

OpenStudy (anonymous):

You didn't. At all. You say it's 0.44 divided by 88? That's not even slightly correct. \[\frac{0.44}{88}=0.005\] Which is not even close. If if you meant the other way around: \[\frac{88}{0.44}=200\] Which is also very wrong. Please stop giving completely wrong answers.

OpenStudy (anonymous):

Umm excuse me? I'm sorry sir...my bad..and why don't you show a little more respect than that too. I'd really appreciate it....and stop posting questions if you don't want help..that's the dumbest thing I have ever heard someone do and also do me a favor and change your attitude towards me and we will be fine and if you can't do that then don't bother me or reply anymore..got that?

OpenStudy (anonymous):

I didn't post the question. I answered it. You purposely posted a completely wrong answer and then asked for a medal for it. I pointed that out. If you don't want to be corrected, don't be wrong.

OpenStudy (anonymous):

I'm not going to argue with someone on a study website.. grow up

OpenStudy (anonymous):

Oh and I'm sorry I didn't know that you are perfect..and I wasn't the only one who was wrong here either..you were wrong to call me out...I don't who you are and I honestly don't even care but you need to learn the right way to tell someone that they are wrong.. Just drop it i'm done..i am on this website to get help on Trigonometry work and not to argue with some little kid that thinks they are perfect. Got any more comebacks?